The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below: Standard Hours Standard Rate Standard Cost Motor tune-up 2.50 $30.00 $75.00 The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 50 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups: Labor rate variance $ 250 F Labor spending variance $ 200 U Required: 1. Determine the number of actual labor-hours spent on tune-ups during the week. 2. Determine the actual hourly rate of pay for tune-ups last week. (Round your answer to 2 decimal places.)
Answer #1

Labor efficiency variance = 250+200 = 450 U

a) Labor efficiency variance = (Standard hour-actual hour)Standard rate

-450 = (50*2.5*30-30X)

30X = 4200

X(Actual hour) = 140 Hour

b) Labor rate variance = (Standard rate-actual rate)actual hour

250 = (30*140-140X)

140X = 3950

X(actual rate) = 28.21